# External Hippocampus: Topological Cognitive Maps for Guiding Large Language Model Reasoning

### v2

@@ Abstract
This paper proposes the External Hippocampus framework, which models language model reasoning from a cognitive dynamics perspective as the flow of information energy in semantic space. Unlike traditional weight-space optimization methods, this framework constructs topological cognitive maps through dimensionality reduction projection, enabling precise navigation and intervention of energy flow at test time while avoiding substantial computational requirements and demonstrating predictable intervention patterns. The method effectively addresses the cognitive deadlock problem in multi-step reasoning for small models. Experiments on models $\leq$7B parameters show: map-guided methods achieve 81.20% accuracy on 500 challenging problems (relative baseline +16.80%), reduce reasoning time by $\geq 15\times$, with key findings revealing that reasoning stagnation manifests as “Cognitive Vortex” and low-entropy potential wells, while temperature perturbations effectively restart energy flow. The framework requires no additional training, possesses autonomous growth capability, and provides an efficient and controllable topological-aware solution for small model reasoning.
Keywords: External Hippocampus; Topological Cognitive Maps; Topological Awareness (Test-time Intervention); Cognitive Deadlock (Low-entropy Attractors); Small Models ($\leq$7B)
1 Introduction
As parameters and training data increase, language models show improved performance on many tasks and exhibit new capabilities. However, performance remains unstable on multi-step tasks requiring continuous, reliable reasoning. Existing research indicates that moderate increases in test-time computation (allowing models to conduct more thorough expansion and checking during reasoning) can improve performance on such tasks ([1, 2, 3, 4]). Methods like Chain-of-Thought (CoT), Tree-of-Thoughts (ToT), and self-consistency suggest that improvements come not only from larger model scale but also from structured organization and timely intervention in the reasoning process.
@@ 1 Introduction
  • Characterize the low-entropy nature of cognitive deadlock and demonstrate that targeted perturbations can alleviate low-entropy attractors ([1, 5, 2]).
  • On two sets of $\leq$7B models using a fixed 500-problem challenging subset, map-guided iterative refinement and majority voting achieve stable improvements over zero-shot baselines; compared to CoT self-consistency, our method is more robust and matches large-scale sampling methods with lower computational cost ([3, 6, 7, 8]).
2 Method
2.1 Experimental Setup and Datasets
Models: We use two small model series (both $\leq$7B parameters) for experiments: (1) Qwen-2.5-3B-Instruct (3.09B parameters, 36 layers, 16 attention heads [9]); (2) Phi-3-mini-4k-instruct (3.8B parameters, 32 layers, 32 attention heads). All experiments run on NVIDIA GPUs using FP16 precision, with caching disabled (use_cache=False) to ensure determinism.
@@ 2.1 Experimental Setup and Datasets
Reproducibility: All experiments use fixed random seed (seed=42) to ensure reproducibility. Cognitive map construction is deterministic (online clustering algorithm), navigator training uses fixed initialization. Experimental code and configuration files are open-sourced, all key parameters are explicitly stated in the paper. The complete reproduction kit will be released upon publication.
@@
2.2 System Architecture
The External Hippocampus system consists of three core components (Figure 1): (1) Cognitive Map Builder, which extracts semantic embeddings from reasoning trajectories and constructs topological graphs; (2) Learned Navigator, which trains an MLP scoring function to predict transition values; (3) Intervention Engine, which dynamically injects hints or perturbations based on map states. The entire system updates online at test time for continuous learning.
2.3 Mathematical Foundation: Cognitive Deadlock and Entropy Dynamics
2.3.1 Cognitive Manifold Hypothesis
@@
2.3.2 Trust Score Dynamics
@@
2.3.3 Entropy Characteristics of Cognitive Deadlock
@@
2.3.4 Perturbation Recovery Mechanism
@@ 2.4 Cognitive Map Construction
We represent the reasoning process as trajectories in semantic latent space ([10, 11, 12]). Given a reasoning step $x_{t}$, we use BAAI/bge-small-en-v1.5 to map it to vector $v_{t}=\text{Embed}(x_{t})$ ([13]). We employ an online nearest-neighbor clustering algorithm to discretize this continuous space into a set of cognitive states $S=\{s_{1},...,s_{N}\}$.
@@
2.5 Learned Navigator Strategy
@@
To guide reasoning, we train a lightweight MLP scorer $f_{\theta}(s_{t},s_{t+1})\to[0,1]$ to predict the value of potential transitions ([14, 15]). This scorer essentially performs “Cognitive Path Value Assessment”. The scorer architecture is a three-layer MLP (input dimension $2d+2$, hidden layer 256 dimensions, output 1 dimension), with input features:
@@
Intervention Strategy: We use a balanced intervention strategy ([1, 2, 3]):
@@
2.6 Map-Guided Deep Thinking
We implemented an iterative refinement loop ([8, 16]). The algorithm flow is as follows:
@@ 2.6 Map-Guided Deep Thinking
2.7 Design Decisions and Empirical Analysis
@@
3 Results
3.1 Topological Anatomy of Reasoning
We visualized cognitive maps for Qwen-2.5-3B and Phi-3-mini constructed from 10k reasoning trajectories ([17, 18, 19, 20, 21]). As shown in Figure 2 and 3, the two models exhibit distinct topological characteristics:
@@ 3.1 Topological Anatomy of Reasoning
3.2 Physics of Cognitive Deadlock
@@
3.3 Intervention Dynamics: Inverted-U Curve
We evaluated different intervention intensities on the fixed 500 hard problem subset for Phi-3-mini and Qwen-2.5-3B (compared with recent prompting/self-consistency/ToT/programmatic reasoning methods [1, 2, 3, 16, 22, 23, 7, 8, 24]). As shown in Figure 5, performance for both models follows an inverted-U curve:
@@ 3.3 Intervention Dynamics: Inverted-U Curve
3.4 Main Performance Comparison
We compared map-guided deep thinking (iterative refinement + majority voting) with strong baselines on the hard subset for Phi-3-mini and Qwen-2.5-3B ([1, 2, 3]). As shown in Figure 6, our method achieved stable improvements on both models.
@@ 3.4 Main Performance Comparison
Table 1: Performance comparison on 500 Hard Reasoning Problems. For Phi-3-mini, the method improves over CoT-SC by +3.4%, suggesting that map-guided direction is more efficient than blind sampling. For Qwen-2.5-3B, CoT-SC degrades performance (-4.60%), while map-guided iterative refinement achieves +16.80%, underscoring the value of topology-aware intervention.
It should be noted that baseline performances reported in Table 2 (e.g., Phi-3-mini 83.00% and Qwen-2.5-3B 64.40% on test_set_500_fixed) correspond to single-turn reasoning on the same “comprehensive hard” 500-problem test set; while the Phi-3-mini (68.0% to 71.4%) and Qwen-2.5-3B (64.40% to 81.20%) discussed in this subsection and iterative refinement experiments are based on a harder subset of 500 problems selected from the same corpus. Therefore, absolute success rates should not be directly compared horizontally across tables. This section focuses on differences in relative improvement and required iteration rounds between different models under comparable difficulty settings (same 500 hard samples): Phi-3-mini achieved limited but stable improvement from a higher baseline, while Qwen-2.5-3B, despite a lower starting point, achieved significant performance leaps through shallow iteration and fine-grained intervention on the same 500 hard samples.
3.5 Online Learning and Growth
Figure 7 demonstrates the system’s ability to learn at test time on Phi-3-mini and Qwen-2.5-3B ([16, 22, 23, 24]). Experimental setup: starting from an empty map, processing 200 samples per round for 5 rounds (total 1000 samples). After each round, the system (1) saves the updated cognitive map, (2) retrains the navigator policy from the map, and (3) loads the new policy for the next round.
@@ 3.5 Online Learning and Growth
[…]
3.6 Cross-Model Navigation: Topology-Dependent Strategy and Diagnostic Value
The distinct topologies of Qwen-2.5-3B and Phi-3-mini require different navigation strategies ([19, 20, 17, 18]). Table 2 summarizes the performance of learned navigator strategies on both models.
@@ 3.6 Cross-Model Navigation: Topology-Dependent Strategy and Diagnostic Value
Table 2: Cross-model navigation performance. The Learned Navigator (bold) is the actual deployed system, using dynamic perturbation and hints rather than complete node masking. For Qwen-2.5-3B, the generic strategy (P=0.5, H=0.6) achieves 64.80% (+0.40%), while a customized strategy (P=0.4, H=0.75) achieves 64.60% (+0.20%), demonstrating that generic parameters are near-optimal. Complete node masking (Ablation) fails (-1.60%), proving the necessity of fine-grained intervention. The key finding is the diagnostic value of cognitive maps: they reveal topology-dependent strategy requirements, with Qwen’s “Tectonic Plates” structure requiring fine-grained negative guidance rather than aggressive masking.
@@
3.7 Cross-Domain Transfer: Semantic Isolation Boundary
To test the generality of cognitive maps, we applied the math reasoning map (trained on 10k math problems) to a code generation task (HumanEval, 164 Python problems) ([25, 26]). Results revealed a strict domain boundary:
@@ 3.7 Cross-Domain Transfer: Semantic Isolation Boundary
Table 3: Zero-shot cross-domain transfer from Math to Code. The identical performance suggests strong semantic isolation under this setup.
@@
A direct implication is: “The shape of thought” is not necessarily universal. Math logic and coding logic may occupy distant regions in the semantic manifold, thus potentially requiring (1) domain-specific maps, (2) cross-domain alignment layers, or (3) multi-modal cognitive architectures capable of bridging semantic gaps ([27, 28, 16]).
4 Discussion
4.1 Generalization Boundary: Different Domains, Different Worlds
While our system performed excellently in math reasoning, cross-domain experiments revealed a clear boundary ([25, 26, 1]). Applying the math-derived cognitive map to code generation (HumanEval) showed no improvement (66.46% vs 66.46%). This negative result is scientifically valuable: it supports that cognitive maps are more likely domain-specific manifolds rather than universal reasoning templates.
Semantic Isolation Hypothesis: Our topological analysis (Section 3.1) revealed that different models exhibit different “signatures” (Qwen’s “Plate Tectonics” vs. Phi-3’s “Neural Archipelago”). This divergence is more pronounced across domains, as Wittgenstein said: “The limits of my language mean the limits of my world.” Math reasoning and code generation occupy fundamentally different regions of the semantic embedding space, possessing mutually incompatible logical topologies. Therefore, a cognitive map (language) constructed based on math context naturally cannot describe or navigate the semantic space (world) of code generation.
@@ 4.1 Generalization Boundary: Different Domains, Different Worlds
    Cross-Domain Alignment: Train a projection layer to map embeddings from one domain to another, achieving “translation” between manifolds ([27, 28]).
@@
This result triggers profound reflection on the nature of “generalization”, which we call “Seed Theory”: Large model capabilities do not emerge from nothing but rely on structural “seeds” buried in pre-training data. Capabilities can only “sprout” when the target task’s semantic structure has topological connectivity with the pre-buried seeds on the manifold. The failure of transfer from math to code actually proves that the so-called “rootless seedling” type of generalization might be a misunderstanding—the significance of piling up parameters lies in expanding the coverage of the “seed bank”, not creating magical capabilities detached from manifold constraints.
4.2 Outlook on Test-time Compute
Compared to strategies relying on massive sampling ([3, 6]), structured memory (maps) serves as an effective enhancement for test-time compute. By deciding where to expand reasoning and intervene, language models achieve more stable performance at lower costs. This “External Hippocampus” approach offers a direction for efficient reasoning enhancement ([29, 30]).
Ultimately, we can view the External Hippocampus as an “Energy Landscape Shaper”. It not only passively records paths but actively fills deep potential wells (via perturbation) and excavates new channels (via hints), reshaping the model’s reasoning topology. This dynamic intervention in the cognitive manifold might be a new path toward more robust artificial intelligence.
5 Conclusion
@@
Copyright Statement
This work is licensed under a Creative Commons Attribution 4.0 International License. You are free to share, copy, distribute and transmit the work, and to adapt the work, for any purpose (including commercial use), as long as you attribute the work to the original authors and provide a link to the license.
License link: https://creativecommons.org/licenses/by/4.0/

### v3

@@ Abstract
This paper proposes the External Hippocampus framework, which models language model reasoning from a cognitive dynamics perspective as the flow of information energy in semantic space. Unlike traditional weight-space optimization methods, this framework constructs topological cognitive maps through dimensionality reduction projection, enabling precise navigation and intervention of energy flow at test time while avoiding substantial computational requirements and demonstrating predictable intervention patterns. The method effectively addresses the cognitive deadlock problem in multi-step reasoning for small models. Experiments on models $\leq$7B parameters show: map-guided methods achieve 81.20% accuracy on 500 challenging problems (relative baseline +16.80%), reduce reasoning time by $\geq 15\times$, with key findings revealing that reasoning stagnation manifests as “Cognitive Vortexes” and low-entropy potential wells, while temperature perturbations effectively restart energy flow. The framework requires no additional training, possesses autonomous growth capability, and provides an efficient and controllable topological-aware solution for small model reasoning.
I Introduction
As parameters and training data increase, language models show improved performance on many tasks and exhibit new capabilities. However, performance remains unstable on multi-step tasks requiring continuous, reliable reasoning. Existing research indicates that moderate increases in test-time computation (allowing models to conduct more thorough expansion and checking during reasoning) can improve performance on such tasks ([wei2022chain, wang2023selfconsistency, yao2023tree, lightman2023let]). Methods like Chain-of-Thought (CoT), Tree-of-Thoughts (ToT), and self-consistency suggest that improvements come not only from larger model scale but also from structured organization and timely intervention in the reasoning process.
@@ I Introduction
  • Characterize the low-entropy nature of cognitive deadlock and demonstrate that targeted perturbations can alleviate low-entropy attractors ([wei2022chain, kojima2022large, wang2023selfconsistency]).
  • On two sets of $\leq$7B models using a fixed 500-problem challenging subset, map-guided iterative refinement and majority voting achieve stable improvements over zero-shot baselines; compared to CoT self-consistency, our method is more robust and matches large-scale sampling methods with lower computational cost ([yao2023tree, zhou2023leasttomost, chen2022react, jiang2023selfrefine]).
II Method
II-A Experimental Setup and Datasets
Models: We use two small model series (both $\leq$7B parameters) for experiments: (1) Qwen-2.5-3B-Instruct (3.09B parameters, 36 layers, 16 attention heads [vaswani2017attention]); (2) Phi-3-mini-4k-instruct (3.8B parameters, 32 layers, 32 attention heads). All experiments run on NVIDIA GPUs using FP16 precision, with caching disabled (use_cache=False) to ensure determinism.
@@ II-A Experimental Setup and Datasets
Reproducibility: All experiments use fixed random seed (seed=42) to ensure reproducibility. Cognitive map construction is deterministic (online clustering algorithm), navigator training uses fixed initialization. Experimental code and configuration files are open-sourced, all key parameters are explicitly stated in the paper. The complete reproduction kit (including source code, pre-trained models, datasets, and run scripts) is provided as supplementary material.
@@
II-B System Architecture
The External Hippocampus system consists of three core components: (1) Cognitive Map Builder, which extracts semantic embeddings from reasoning trajectories and constructs topological graphs; (2) Learned Navigator, which trains an MLP scoring function to predict transition values; (3) Intervention Engine, which dynamically injects hints or perturbations based on map states. The entire system updates online at test time for continuous learning.
II-C Mathematical Foundation: Cognitive Deadlock and Entropy Dynamics
II-C1 Cognitive Manifold Hypothesis
@@
II-C2 Trust Score Dynamics
@@
II-C3 Entropy Characteristics of Cognitive Deadlock
@@
II-C4 Perturbation Recovery Mechanism
@@ II-D Cognitive Map Construction
We represent the reasoning process as trajectories in semantic latent space ([reimers2019sentence, gao2021retrieval, karpukhin2020dense]). Given a reasoning step $x_{t}$, we use BAAI/bge-small-en-v1.5 to map it to vector $v_{t}=\text{Embed}(x_{t})$ ([xiao2023bge]). We employ an online nearest-neighbor clustering algorithm to discretize this continuous space into a set of cognitive states $S=\{s_{1},...,s_{N}\}$.
@@
II-E Learned Navigator Strategy
@@
To guide reasoning, we train a lightweight MLP scorer $f_{\theta}(s_{t},s_{t+1})\to[0,1]$ to predict the value of potential transitions ([rumelhart1986learning, lecun2015deep]). This scorer essentially performs “Cognitive Path Value Assessment”. The scorer architecture is a three-layer MLP (input dimension $2d+2$, hidden layer 256 dimensions, output 1 dimension), with input features:
@@
Intervention Strategy: We use a balanced intervention strategy ([wei2022chain, wang2023selfconsistency, yao2023tree]):
@@
II-F Map-Guided Deep Thinking
We implemented an iterative refinement loop ([jiang2023selfrefine, chen2023program]). The algorithm flow is as follows:
@@ II-F Map-Guided Deep Thinking
II-G Design Decisions and Empirical Analysis
@@
III Results
III-A Topological Anatomy of Reasoning
We visualized cognitive maps for Qwen-2.5-3B and Phi-3-mini constructed from 10k reasoning trajectories ([microsoft2024phi3, qwen2024qwen2_5, besta2024topology]). As shown in Figure 2 and 3, the two models exhibit distinct topological characteristics:
@@ III-A Topological Anatomy of Reasoning
III-B Physics of Cognitive Deadlock
@@
III-C Intervention Dynamics: Inverted-U Curve
We evaluated different intervention intensities on the fixed 500 hard problem subset for Phi-3-mini and Qwen-2.5-3B (compared with recent prompting/self-consistency/ToT/programmatic reasoning methods [wei2022chain, wang2023selfconsistency, yao2023tree, chen2023program, gao2023pal, madaan2023selfverify, chen2022react, jiang2023selfrefine, shinn2023reflexion]). As shown in Figure 5, performance for both models follows an inverted-U curve:
@@ III-C Intervention Dynamics: Inverted-U Curve
TABLE I: Performance comparison on 500 Hard Reasoning Problems. For Phi-3-mini, the method improves over CoT-SC by +3.4%, suggesting that map-guided direction is more efficient than blind sampling. For Qwen-2.5-3B, CoT-SC degrades performance (-4.60%), while map-guided iterative refinement achieves +16.80%, underscoring the value of topology-aware intervention.
III-D Main Performance Comparison
We compared map-guided deep thinking (iterative refinement + majority voting) with strong baselines on the hard subset for Phi-3-mini and Qwen-2.5-3B ([wei2022chain, wang2023selfconsistency, yao2023tree]). As shown in Figure 6, our method achieved stable improvements on both models.
It should be noted that baseline performances reported in Table II (e.g., Phi-3-mini 83.00% and Qwen-2.5-3B 64.40% on test_set_500_fixed) correspond to single-turn reasoning on the same “comprehensive hard” 500-problem test set; while the Phi-3-mini (68.0% to 71.4%) and Qwen-2.5-3B (64.40% to 81.20%) discussed in this subsection and iterative refinement experiments are based on a harder subset of 500 problems selected from the same corpus. Therefore, absolute success rates should not be directly compared horizontally across tables. This section focuses on differences in relative improvement and required iteration rounds between different models under comparable difficulty settings (same 500 hard samples): Phi-3-mini achieved limited but stable improvement from a higher baseline, while Qwen-2.5-3B, despite a lower starting point, achieved significant performance leaps through shallow iteration and fine-grained intervention on the same 500 hard samples.
[…]
III-E Online Learning and Growth
Figure 7 demonstrates the system’s ability to learn at test time on Phi-3-mini and Qwen-2.5-3B ([chen2023program, gao2023pal, madaan2023selfverify, shinn2023reflexion]). Experimental setup: starting from an empty map, processing 200 samples per round for 5 rounds (total 1000 samples). After each round, the system (1) saves the updated cognitive map, (2) retrains the navigator policy from the map, and (3) loads the new policy for the next round.
@@ III-E Online Learning and Growth
TABLE II: Cross-model navigation performance. The Learned Navigator (bold) is the actual deployed system, using dynamic perturbation and hints rather than complete node masking. For Qwen-2.5-3B, the generic strategy (P=0.5, H=0.6) achieves 64.80% (+0.40%), while a customized strategy (P=0.4, H=0.75) achieves 64.60% (+0.20%), demonstrating that generic parameters are near-optimal. Complete node masking (Ablation) fails (-1.60%), proving the necessity of fine-grained intervention. The key finding is the diagnostic value of cognitive maps: they reveal topology-dependent strategy requirements, with Qwen’s “Tectonic Plates” structure requiring fine-grained negative guidance rather than aggressive masking.
III-F Cross-Model Navigation: Topology-Dependent Strategy and Diagnostic Value
The distinct topologies of Qwen-2.5-3B and Phi-3-mini require different navigation strategies ([microsoft2024phi3, qwen2024qwen2_5, touvron2023llama, dubey2023llama2]). Table II summarizes the performance of learned navigator strategies on both models.
@@ III-F Cross-Model Navigation: Topology-Dependent Strategy and Diagnostic Value
III-G Cross-Domain Transfer: Semantic Isolation Boundary
To test the generality of cognitive maps, we applied the math reasoning map (trained on 10k math problems) to a code generation task (HumanEval, 164 Python problems) ([chen2021evaluating, hendrycks2021measuring]). Results revealed a strict domain boundary:
@@ III-G Cross-Domain Transfer: Semantic Isolation Boundary
TABLE III: Zero-shot cross-domain transfer from Math to Code. The identical performance suggests strong semantic isolation under this setup.
@@
A direct implication is: “The shape of thought” is not necessarily universal. Math logic and coding logic may occupy distant regions in the semantic manifold, thus potentially requiring (1) domain-specific maps, (2) cross-domain alignment layers, or (3) multi-modal cognitive architectures capable of bridging semantic gaps ([radford2021learning, liu2023visual, chen2023program]).
IV Discussion
IV-A Generalization Boundary: Different Domains, Different Worlds
While our system performed excellently in math reasoning, cross-domain experiments revealed a clear boundary ([chen2021evaluating, hendrycks2021measuring, wei2022chain]). Applying the math-derived cognitive map to code generation (HumanEval) showed no improvement (66.46% vs 66.46%). This negative result is scientifically valuable: it supports that cognitive maps are more likely domain-specific manifolds rather than universal reasoning templates.
Semantic Isolation Hypothesis: Our topological analysis (Section III-A) revealed that different models exhibit different “signatures” (Qwen’s “Plate Tectonics” vs. Phi-3’s “Neural Archipelago”). This divergence is more pronounced across domains, as Wittgenstein said: “The limits of my language mean the limits of my world.” Math reasoning and code generation occupy fundamentally different regions of the semantic embedding space, possessing mutually incompatible logical topologies. Therefore, a cognitive map (language) constructed based on math context naturally cannot describe or navigate the semantic space (world) of code generation.
@@ IV-A Generalization Boundary: Different Domains, Different Worlds
    Cross-Domain Alignment: Train a projection layer to map embeddings from one domain to another, achieving “translation” between manifolds ([radford2021learning, liu2023visual]).
@@
This result triggers reflection on the nature of “generalization”, which we call “Seed Theory”: Large model capabilities do not emerge from nothing but rely on structural “seeds” buried in pre-training data. Capabilities can only “sprout” when the target task’s semantic structure has topological connectivity with the pre-buried seeds on the manifold. The failure of transfer from math to code actually proves that the so-called “rootless seedling” type of generalization might be a misunderstanding—the significance of piling up parameters lies in expanding the coverage of the “seed bank”, not creating capabilities detached from manifold constraints.
IV-B Outlook on Test-time Compute
Compared to strategies relying on massive sampling ([yao2023tree, zhou2023leasttomost]), structured memory (maps) serves as an effective enhancement for test-time compute. By deciding where to expand reasoning and intervene, language models achieve more stable performance at lower costs. This “External Hippocampus” approach offers a direction for efficient reasoning enhancement ([lecun2022pathways, bubeck2023sparks]).
Ultimately, we can view the External Hippocampus as an “Energy Landscape Shaper”. It not only records paths but actively fills deep potential wells (via perturbation) and excavates new channels (via hints), reshaping the model’s reasoning topology. This dynamic intervention in the cognitive manifold suggests a new path toward more robust artificial intelligence.
V Conclusion